### v2

# Stabbing non-piercing sets and face lengths in large girth plane graphs

Abstract
We show that a non-piercing family of connected planar sets with bounded independence number can be stabbed with a constant number of points. As a consequence, we answer a question of Axenovich, Kießle and Sagdeev about the largest possible face length of an edge-maximal plane graph with girth at least $\ell$.
1 Introduction
Define a region as a connected planar compact set whose boundary consists of a finite number of disjoint Jordan curves. One of these curves is the outer boundary of the region, while the rest cut out “holes”. A family $\mathcal{F}$ of regions is in general position if for any two regions from $\mathcal{F}$ their boundaries intersect in finitely many points. [^1] This is a technical condition that was introduced in [19] and makes many of the arguments simpler, though most often could be omitted. We need to assume it as we will use a result from [19] which was proved under this assumption, although we might get rid of it due to a very recent result [21] which still needs to be verified. Such a family $\mathcal{F}$ is non-piercing if $F\setminus G$ is connected for any two regions $F,G\in\mathcal{F}$. For example, the family of all disks is non-piercing and, more generally, so is a pseudo-disk family, defined as a family of simply connected regions whose boundaries intersect pairwise at most twice; these include families formed by homothetic [^2] A homothetic copy is a copy that is translated and scaled by a positive scalar factor. copies of a fixed convex set. However, a family of axis-parallel rectangles in general position is not necessarily non-piercing, as two rectangles can cross each other without any of them containing a vertex of the other.
For a family $\mathcal{F}$, the independence number $\nu(\mathcal{F})$ is the size of the largest subfamily of pairwise disjoint sets, that is, the smallest number such that among any $\nu(\mathcal{F})+1$ sets there are two that intersect. The piercing number $\tau(\mathcal{F})$ is the least number of points that pierce $\mathcal{F}$, that is, the size of the smallest point set that intersects every set from $\mathcal{F}$. Our main result is the following.
Theorem 1.
There is a function $f$ such that if $\mathcal{F}$ is a family of non-piercing regions, then $\tau(\mathcal{F})\leq f(\nu(\mathcal{F}))$.
Note that this implies that the disjointness graph of a family of non-piercing regions is $\chi$-bounded. [^3] For the definition and a survey of $\chi$-boundedness, see [24]. For results similar to Theorem 1 about homothetic copies of a fixed convex set with explicit bounds, see [12].
In particular, if $\mathcal{F}$ is a family of pairwise intersecting, non-piercing regions, then there exists an absolute constant $T$ such that $\mathcal{F}$ can be pierced with at most $T$ points. This result was proved earlier for pseudo-disk families [1] with different methods, using sweepings, which do not generalize to non-piercing families. Instead, our proof, which can be found in Section 2, uses the standard machinery developed to prove $(p,q)$-theorems. We do not know the best possible value for $T$. In case of pairwise intersecting disks, this is a well-studied problem where we know that the optimal value is four, for which by now there are several different proofs, see [9, 25, 6]. It is entirely possible that the answer in case of pairwise intersecting non-piercing regions is also four but at the moment this is not known even for pseudo-disks.
As an application of Theorem 1, we answer a recent question of Axenovich, Kießle and Sagdeev [3] about the largest possible face length of an edge-maximal plane graph with girth at least $\ell$, which was our main motivation.
Theorem 2.
Suppose that $G$ is a plane graph with girth at least $\ell$, and that $G$ is edge-maximal with regards to these two properties. Then the length of any facial cycle of $G$ is at most $K\ell$ for some absolute constant $K$.
The exact statement of the problem and the proof can be found in Section 3.
In Section 4, we show a generalization of Theorem 2 for graphs drawn on other orientable surfaces:
{restatable}
Thfacesgen For every genus $g$, there exists a constant $c_{g}$ such that for every edge-maximal graph $G$ with girth $\ell$ drawn on an orientable surface with genus $g$, the maximal face length of $G$ is bounded by $c_{g}\cdot\ell$.
Remarks.
Since the first version of our paper appeared on arXiv, Theorem 1 was improved by Huang, Qi, Rong and Xu [15], who showed that $\tau(\mathcal{F})\leq c\nu(\mathcal{F})^{9}$ for some constant $c$. They used the notion of dyadic VC-dimension and an old result of Ding, Seymour, and Winkler [11] to obtain better bounds.
Later the bound was further improved to the optimal $\tau(\mathcal{F})=O(\nu(\mathcal{F}))$ by Keller and Smorodinsky [16] who used a recent result of Dalal, Gangopadhyay, Raman, and Ray [8] that generalized a result of Pinchasi [18] from pseudo-disks to non-piercing regions. This result is that in every family $\mathcal{F}$ of non-piercing regions there is an $F\in\mathcal{F}$ which is “small” in the sense that there cannot be $157$ pairwise disjoint members of $\mathcal{F}$ that intersect $F$.
In the new version of the paper [3] (version 2 on arXiv), Axenovich, Kießle, Sagdeev, and Zhukovskii also proved their original question, giving a concrete upper bound of $8\ell-13$, thus improving the bound of our Theorem 2. They also raised the problem for surfaces of higher genus $g$, for which they first obtained a polynomial, later an $O_{g}(\ell^{2})$ bound. This inspired us to include Theorem 2, which improves this to $O_{g}(\ell)$. In the third version of their paper [3] (version 3 on arXiv), they also proved a bound of $O_{\ell}(g)$. They also raised the problem whether a bound of $O(g\ell)$ holds.
2 Proof of Theorem 1
In this section we present the proof of Theorem 1. We start with some definitions, related to $(p,q)$-theorems; for a complete survey of such results, see [14].
Following Hadwiger and Debrunner [13], we say that a family $\mathcal{G}$ has the $(p,q)$-property, if for every subfamily $\mathcal{G}\subset\mathcal{F}$ with $|\mathcal{G}|=p$, there exists a subsubfamily $\mathcal{H}\subset\mathcal{G}$ of size $q$ with a non-empty intersection $\cap\mathcal{H}\neq\emptyset$. In other words, from every $p$ sets from $\mathcal{F}$, some $q$ intersect. It was shown by Alon and Kleitman [2] that if a family $\mathcal{F}$ of compact convex sets in $\mathbb{R}^{d}$ satisfies the $(p,q)$-property for any $p\geq q\geq d+1$, then $\mathcal{F}$ can be pierced with $T(p,q)$ points. Later, this result was extended from convex sets to several other families. For the version that we need, we need to define the Vapnik-Chervonenkis dimension. The VC-dimension of a family $\mathcal{F}$ is the largest $d$ for which exists a set of $d$ elements, $X$, such that for every subset $Y\subset X$ there exists a set $F\subset\mathcal{F}$ such that $X\cap F=Y$. The dual VC-dimension $d^{*}$ of $\mathcal{F}$ is the VC-dimension of the dual family $\mathcal{F}^{*}$, in which the roles of elements and sets are swapped, with the containment relation reversed. It is well-known and easy to see that $d^{*}\leq 2^{d}$. Matoušek [17] showed that bounded VC-dimension and an appropriate $(p,q)$-property imply the existence of a small hitting set.
Theorem 3 (Matoušek [17]).
If the dual VC-dimension of $\mathcal{F}$ is at most q1q-1, and $\mathcal{F}$ satisfies the $(p,q)$-property for some $p\geq q$, then the sets of $\mathcal{F}$ can be hit with at most $T$ points, where $T$ is a constant depending on $p$ and $q$.
Therefore, in order to prove Theorem 1, it would be sufficient to show that non-piercing regions have bounded (dual) VC-dimension, and if a family $\mathcal{F}$ of non-piercing regions has bounded independence number $\nu(\mathcal{F})$, then they also satisfy the $(p,q)$-property for some large enough $q$.
Lemma 4.
If $\mathcal{F}$ is a non-piercing family of regions, then the VC-dimension and the dual VC-dimension of $\mathcal{F}$ are at most $4$.
Our proof is somewhat similar to [5] where pseudo-disks were considered.
Proof.
Suppose for a contradiction that there exists a planar point set $\{x_{1},\ldots,x_{5}\}$ shattered by $\mathcal{F}$. This implies that for each pair of points $x_{i},x_{j}$, there exists $F_{i,j}\in\mathcal{F}$ which contains $x_{i}$ and $x_{j}$ but does not contain any of the other three points. Since $F_{i,j}$ is connected, there exists a simple curve $\gamma_{i,j}\subset F_{i,j}$ whose ends are $x_{i}$ and $x_{j}$. Let us fix one such curve for each of the ten pairs $(i,j)$ such that any two curves intersect a finite number of times. These ten curves form a planar embedding of the complete graph $K_{5}$ with vertices $\{x_{1},\ldots,x_{5}\}$, therefore, by the strong Hanani-Tutte theorem [26], there are two independent edges, without loss of generality, $\gamma_{1,2}$ and $\gamma_{3,4}$, that cross an odd number of times. As $F_{1,2}\setminus F_{3,4}$ is connected because of the non-piercing property, we can extend $\gamma_{1,2}$ to a Jordan-curve $\bar{\gamma}_{1,2}\subset F_{1,2}$ that still crosses $\gamma_{3,4}\subset F_{3,4}$ an odd number of times. Similarly, we can extend $\gamma_{3,4}$ to a Jordan-curve $\bar{\gamma}_{3,4}\subset F_{3,4}$ that still crosses $\bar{\gamma}_{1,2}$ an odd number of times. But then these two closed curves would violate the Jordan curve theorem.
The argument for the dual VC-dimension is similar. Take five sets, $F_{1},\ldots,F_{5}\in\mathcal{F}$ that are shattered. Take five points, $\{x_{1},\ldots,x_{5}\}$, such that $x_{i}\in F_{i}\setminus\cup_{k\neq i}F_{k}$ for each $i$ and ten points such that $y_{i,j}\in(F_{i}\cap F_{j})\setminus\cup_{k\neq i,j}F_{k}$ for each $i<j$. Take simple curves $\gamma_{i,j}\subset F_{i}$ whose ends are $x_{i}$ and $x_{i,j}$ and $\gamma_{j,i}\subset F_{j}$ whose ends are $x_{j}$ and $x_{i,j}$ for each $i<j$. The ten concatenations $\gamma_{i,j}\gamma_{j,i}$ of the curves form a planar embedding of the complete graph $K_{5}$ with vertices $\{x_{1},\ldots,x_{5}\}$, therefore, by the strong Hanani-Tutte theorem, there are two independent edges that cross an odd number of times. As each edge is the concatenation of two curves, $\gamma_{i,j}$ and $\gamma_{j,i}$, there are also two curves, without loss of generality, $\gamma_{1,2}\subset F_{1}$ and $\gamma_{3,4}\subset F_{3}$, that cross an odd number of times. From here the proof is the same as before. As $F_{1}\setminus F_{3}$ is connected because of the non-piercing property, we can extend $\gamma_{1,2}$ to a Jordan-curve $\bar{\gamma}_{1,2}\subset F_{1}$ that still crosses $\gamma_{3,4}\subset F_{3}$ an odd number of times. Similarly, we can extend $\gamma_{3,4}$ to a Jordan-curve $\bar{\gamma}_{3,4}\subset F_{3}$ that still crosses $\bar{\gamma}_{1,2}$ an odd number of times. But then these two closed curves would violate the Jordan curve theorem. ∎
Now we only need to show that a $(p,q)$-property holds for some $p\geq q\geq 5$ in every family $\mathcal{F}$ of non-piercing regions with bounded independence number $\nu(\mathcal{F})$. We first make some definitions.
For a family $\mathcal{F}$ and collection of elements $P$, define the dual intersection hypergraph $\mathcal{H}(\mathcal{F},P)$ such that the vertices correspond to members of $\mathcal{F}$, while hyperedges correspond to elements of $P$ such that for every $p\in P$ the vertex set $H_{p}=\{F\in\mathcal{F}:p\in F\}$ forms a hyperedge. The Delaunay graph $\mathcal{D(\mathcal{F})}$ is the subgraph of $\mathcal{H}(\mathcal{F},P)$ that contains only the hyperedges with exactly two vertices, that is, a pair of vertices corresponding to the sets $F,G\in\mathcal{F}$ are connected by an edge if there is an element $p\in F\cap G$ which is not contained in any other member of $\mathcal{F}$. Raman and Ray proved (in a much more general form) that if $\mathcal{F}$ is a family of non-piercing regions, then $\mathcal{D(\mathcal{F})}$ is planar.
Corollary 5 (of Raman and Ray [19]).
If $\mathcal{F}$ is a family of non-piercing regions, then the Delaunay graph $\mathcal{D}(\mathcal{F})$ is planar, therefore, it can have at most $3|\mathcal{F}|$ edges.
Now we are ready to state the last lemma needed to complete the proof.
Lemma 6.
If $\mathcal{F}$ is a family of non-piercing regions with the $(\nu+1,2)$-property, then $\mathcal{F}$ also has the $(p,q)$-property for any $p>3e(\nu+1)\nu q+1$ for any $q\geq 2$, where $e=2.71$... is Euler’s number.
Proof.
The proof uses the so-called Clarkson-Shor method [7]. Fix some $q\geq 2$ and $p>3e(\nu+1)\nu q+1$, and suppose for a contradiction that there exists a subfamily $\mathcal{G}\subset\mathcal{F}$ of size $p$ without a common intersection. Delete each set from $\mathcal{G}$ with probability $1-\frac{1}{q}$ to obtain the subsubfamily $\mathcal{H}$. The average number of remaining sets is $\frac{p}{q}$, so in expectation the Delaunay graph $\mathcal{D(\mathcal{H})}$ has this many vertices. The probability that two different intersecting sets of $\mathcal{G}$ span an edge of $\mathcal{D(\mathcal{H})}$ is at least $\frac{1}{q^{2}}\left(1-\frac{1}{q}\right)^{q-3}>\frac{1}{e}\frac{1}{q^{2}}$, as any point in the intersection of two sets is contained in at most q3q-3 other sets; if these are all deleted, then we get a Delaunay edge. As from any $\nu+1$ sets in $\mathcal{G}$, there are two that intersect, the number of intersecting set pairs in $\mathcal{G}$ is at least $\frac{\binom{p}{\nu+1}}{\binom{p-2}{\nu-1}}=\frac{p(p-1)}{(\nu+1)\nu}$. Consequently, the expected number of edges is at least $\frac{1}{e}\frac{p(p-1)}{(\nu+1)\nu q^{2}}$. As $\mathcal{G}$ and all its subsystems are non-piercing, by Corollary 5 the expected number of edges after deletion can be at most three times the expected number of vertices, that is, $\frac{1}{e}\frac{p(p-1)}{(\nu+1)\nu q^{2}}\leq 3\frac{p}{q}$, which contradicts $p>3e(\nu+1)\nu q+1$. ∎
Lemmas 4 and 6 imply that $\mathcal{F}$ satisfies the assumptions of Theorem 3 with $q=5$ and some large enough $p$, which implies that $\mathcal{F}$ can be stabbed with a constant number of points. This finishes the proof of Theorem 1. ∎
3 Proof of Theorem 2
In this section, we present the exact statement and the proof of Theorem 2.
First, we introduce the definitions and notation following Axenovich, Kießle and Sagdeev [3]. A plane graph is a graph that is embedded in the plane without crossing edges. A 2-connected plane graph $G$ is $C_{<\ell}$-free if it contains no cycle of length smaller than $\ell$. $G$ is a maximal $C_{<\ell}$-free plane graph if adding any new edge would either create a crossing or a cycle of length less than $\ell$. Define $f_{max}(\ell)$ as the largest possible face length of a 2-connected maximal $C_{<\ell}$-free plane graph.
Axenovich, Kießle and Sagdeev [3] showed that $f_{max}(\ell)=2\ell-3$ for $3\leq\ell\leq 6$ using a former result of Axenovich, Ueckerdt, and Weiner [4]. For larger values of $\ell$, they showed a lower bound of $3\ell-9$ for $7\leq\ell\leq 9$, and $3\ell-12$ for $\ell\geq 10$.
They also showed an upper bound of $2(\ell-2)^{2}+1$ for any $\ell\geq 7$, and asked whether it could be improved to a linear upper bound. We give an affirmative answer to this question.
Theorem 7 (Theorem 2, restated).
$f_{max}(\ell)\leq K\ell$ for some absolute constant $K$.
We start with a simple observation made by Axenovich, Kießle and Sagdeev [3]. For two vertices $u,v$, let $d(u,v)$ denote their distance in $G$, and define the distance of a vertex $u$ and an edge $e=vw$ as $d(u,e)=\max\{d(u,v),d(u,w)\}$.
Observation 8 ([3]).
For any two vertices $u,v$ of a facial cycle of a maximal $C_{<\ell}$-free graph, their distance $d(u,v)\leq\ell-2$.
Proof.
If $d(u,v)\geq\ell-1$, then $u$ and $v$ are non-adjacent, and the edge $uv$ could be added inside the cycle, preserving planarity, contradicting the maximality of our graph. ∎
We first show that it is sufficient to prove Theorem 7 for the case when $\ell$ is even. Suppose that $\ell$ is odd, and consider a plane graph $G$ with girth at least $\ell$ which is maximal with these constraints and has a face with boundary $C$ of length $m$. Subdivide each edge $e=uv$ with a middle vertex $e^{\prime}$, thus creating a graph $G^{\prime}$. It is easy to see that $G^{\prime}$ is planar, and has girth at least $2\ell$. Additionally, it is maximal to these parameters: By Observation 8, any two vertices on a face of $G$ are connected by a path of length at most $\ell-2$ (not necessarily along the face), and every vertex of $G^{\prime}$ either corresponds to a vertex of $G$ or is adjacent to one, therefore, any two vertices on a face of $G^{\prime}$ are connected by a path of length at most $1+2(\ell-2)+1$. This implies that adding an edge between two vertices of $G^{\prime}$ would either violate planarity or create a cycle of length at most $2\ell-1$. Since $G^{\prime}$ has a face of size $2m$, this implies that $2m\leq K\cdot 2\ell$, which implies our bound for $G$ as well.
Fix a maximal $\mathcal{C}_{\ell}$-free plane graph $G$ for some even $\ell$ and a facial cycle $C$ of $G$, we will bound the length $m$ of $C$ in terms of $\ell$. By Fáry’s theorem, we can assume that $G$ is geometric, that is, each of its edges is a segment, without changing the topology of the embedding.
Let $\bar{B}(x,r)$ denote the closed (euclidean) disk of radius $r$ around a point $x$ in the plane, and for a simple curve $\gamma$, let $\bar{B}(\gamma,r)=\{x\in\mathbb{R}^{2}:\exists p\in\gamma\,\,|x-p|\leq r\}$, the set of points at euclidean distance at most $r$ from (a point of) $\gamma$.
For each $v_{i}\in C$, we define a blow-up of its $(\ell/2-1)$-neighborhood in $G$.
Let $\rho$ be a small positive number that is less than half of the minimum euclidean distance of a vertex and a non-incident edge, and is also less than half of the minimum euclidean distance of two vertices. Additionally, let $\delta<\rho$ be a small enough positive number such that for any two edges $e,f$ incident to a common vertex $u$, the intersection of $\bar{B}(e,\delta)$ and $\bar{B}(f,\delta)$ is contained in $\bar{B}(u,\rho)$. We also define $0<\varepsilon_{1}<\varepsilon_{2}<\dots<\varepsilon_{m}<\frac{\delta}{\ell^{2}}$ to be arbitrary positive numbers.
Finally, for $u\in V(G)$ and $i\in[m]$, let $\bar{B}^{i}(u)=\bar{B}\left(u,\frac{\rho}{d(v_{i},u)+1}+\varepsilon_{i}\right)$, and for $e\in E(G)$ and $i\in[m]$, let $\bar{B}^{i}(e)=\bar{B}\left(e,\frac{\delta}{d(v_{i},e)+1}+\varepsilon_{i}\right)$.
Now we can define the neighborhood regions $N_{i}$ for each vertex $v_{i}$ (see Figure 1).
$N_{i}=\bigcup\limits_{\begin{subarray}{c}u\in V(G)\\ d(v_{i},u)\leq\ell/2-1\end{subarray}}\bar{B}^{i}(u)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cup\bigcup\limits_{\begin{subarray}{c}e\in E(G)\\ d(v_{i},e)\leq\ell/2-1\end{subarray}}\bar{B}^{i}(e).$
$v_{i}$
Figure 1: Illustration for a neighborhood $N_{i}$.
Note that the boundary of any two members of $\mathcal{N}=\{N_{1},N_{2},\dots,N_{m}\}$ intersect in finitely many points as the radii of their blow-ups are perturbed with small amounts, so $\mathcal{N}$ is a family of regions in general position. In order to apply Theorem 1 to $\mathcal{N}$, we need to verify that it satisfies the properties required by the theorem.
Proposition 9.
$N_{i}$ is simply connected for each $i$.
Proof.
Since the girth of $G$ is at least $\ell$, the vertices whose distance from $v_{i}$ is at most $\ell/2-1$ form a tree. ∎
Lemma 10.
The set-system $\mathcal{N}$ is non-piercing.
While quite straight-forward, unfortunately our proof is quite tedious and technical.
Proof.
Suppose for a contradiction that $N_{i}$ pierces $N_{j}$ for some $i\neq j$. Each point in $N_{j}\setminus N_{i}$ is contained in the neighborhood of a vertex or an edge which is closer to, or equally close to $v_{j}$ than to $v_{i}$. Let $v_{j}^{\prime}$ be an arbitrary point in $\bar{B}^{j}(v_{j})\setminus N_{i}$. We will show that any point $x\in N_{j}\setminus N_{i}$ is in the same connected component of $N_{j}\setminus N_{i}$ as $v_{j}^{\prime}$, which contradicts our assumption. Let $u_{x}$ be the vertex for which either $x\in\bar{B}^{j}(u_{x})$, or $x\in\bar{B}^{j}(e)$ for some edge $e$ incident to $u_{x}$ for which $d(v_{j},e)=d(v_{j},u_{x})+1$. We will only discuss the case $\varepsilon_{j}<\varepsilon_{i}$, as the other case goes analogously.
Define $P_{x}=\{u_{0}=v_{j},u_{1},\dots,u_{k}=u_{x}\}$ to be the shortest path from $v_{j}$ to $u_{x}$ in $G$. (This path is unique because $d(u_{x},v_{j})\leq\ell/2-1$.) Note that for each vertex $w\in P_{x}$, $d(w,v_{j})<d(w,v_{i})$, where the inequality is strict because we assumed $\varepsilon_{j}<\varepsilon_{i}$. For an illustration, see Figure 2.
$N_{j}\setminus N_{i}$$N_{i}\setminus N_{j}$$N_{j}\cap N_{i}$$v_{j}$$u_{k-1}$$u_{k}$
Figure 2: An illustration for intersections between different neighborhoods $N_{i}$ and $N_{j}$.
First, we show that points in the neighborhood of the same vertex are in the same connected component of $N_{j}\setminus N_{i}$, though to connect them we might need to leave the neighborhood of the respective vertex.
Proposition 11.
For any $u\in N_{j}$, the set $\bar{B}^{j}\left(u\right)\setminus N_{i}$ is contained in a connected component of $N_{j}\setminus N_{i}$.
Proof.
We may assume $d(u,v_{j})<d(u,v_{i})$, as otherwise $\bar{B}^{j}\left(u\right)\setminus N_{i}$ is empty.
We will use induction on $d(u,v_{i})$.
If $d(u,v_{i})\geq\ell/2$, then $u\notin N_{i}$, so we are done.
To prove the induction step, we need to show that any two points $p_{1},p_{2}$ from $\bar{B}^{j}\left(u\right)\setminus N_{i}$ can be connected in $N_{j}\setminus N_{i}$. We may assume that $p_{1}$ and $p_{2}$ are on the boundary of $\bar{B}^{j}\left(u\right)$, as the segments connecting them to their respective projections from $u$ to the boundary are contained in $N_{j}\setminus N_{i}$.
We claim that there is at most one edge $e$ incident to $u$ for which the other endpoint $u_{e}$ of $e$ satisfies $d(u_{e},v_{i})\leq d(u,v_{i})$. Consider such an edge $e$. In case of equality, there are two adjacent vertices in $G$ both at distance at most $\ell/2-1$ from $v_{i}$. The union of the two shortest paths and the edge between these two vertices creates a non-trivial closed walk of length at most $\ell/2+\ell/2-1$, which is a contradiction. Therefore, for each such edge, there is a shortest path going from $u$ to $v_{i}$ through $e$. If there were two such edges, then their union would create a non-trivial closed walk of length at most $\ell/2-1+\ell/2-1<\ell$, which is again a contradiction.
The points $p_{1}$ and $p_{2}$ divide the boundary circle $\partial\bar{B}^{j}\left(u\right)$ into two arcs. The above claim implies that for one of them, each edge $e$ crossing it satisfies $d(u,v_{i})<d(u_{e},v_{i})$.
This arc is divided into subarcs by edges that cross it. We will show that each pair of subsequent subarcs are in the same connected component of $N_{j}\setminus N_{i}$. This will finish the proof of Proposition 11.
For each crossing edge $e$, both ends of $e$ are closer to $v_{j}$ than to $v_{i}$. This implies that the two boundary segments $\partial\bar{B}^{j}\left(e\right)\setminus\left(\bar{B}^{j}\left(u\right)\cup\bar{B}^{j}\left(u_{e}\right)\right)$ are contained in $N_{j}\setminus N_{i}$, and they connect subsequent arcs to the boundary circle of $\bar{B}^{j}\left(u_{e}\right)$. As $d(u,v_{i})<d(u_{e},v_{i})$, by induction, points on this boundary circle are in the same connected component of $N_{j}\setminus N_{i}$, and so the same is true for points of the subsequent arcs on $\partial\bar{B}^{j}\left(u\right)$. ∎
Now we return to the proof of Lemma 10. We will proceed by induction on the length of the path $P_{x}$, which is $k=d(u_{k},v_{j})$.
The $k=0$ case is implied by Proposition 11.
To prove the induction step, it is sufficient to show that there exist two points $y_{1}\in\bar{B}^{j}\left(u_{k}\right)\setminus N_{i}$ and $y_{2}\in\bar{B}^{j}\left(u_{k-1}\right)\setminus N_{i}$ in the same connected component of $N_{j}\setminus N_{i}$, as Proposition 11 implies that points within the same disk are in the same connected component, and by induction, $y_{2}$ is in the same connected component as $v_{j}^{\prime}$.
Choose $y_{1}$ and $y_{2}$ from the neighborhood of the edge $e=u_{k-1}u_{k}$ such that $y_{1}\in\partial\bar{B}^{j}\left(u_{k}\right)\cap\partial\bar{B}^{j}\left(e\right)$, $y_{2}\in\partial\bar{B}^{j}\left(u_{k-1}\right)\cap\partial\bar{B}^{j}\left(e\right)$ and the segment $y_{1}y_{2}$ is on the boundary $\partial\bar{B}^{j}\left(e\right)$; see Figure 3. This way, the segment $y_{1}y_{2}\subset N_{j}\setminus N_{i}$ which proves that they are in the same connected component of $N_{j}\setminus N_{i}$.
$u_{k}$$u_{k-1}$$y_{2}$$y_{1}$
Figure 3: Two points connected by a segment in $N_{j}\setminus N_{i}$.
This finishes the proof that $x$ and $v_{j}^{\prime}$ are in the same component of $N_{j}\setminus N_{i}$ in the $i<j$ case. The $i>j$ case of Lemma 10 goes very similarly, so we omit the proof. ∎
Lemma 12.
The set-system $\mathcal{N}$ has the $(2,2)$-property, i.e., it is pairwise intersecting.
Proof.
Take two arbitrary vertices $v_{i},v_{j}\in C$. By Observation 8, $d(v_{i},v_{j})\leq\ell-2$. Since $\ell$ is even, this implies that there is a vertex $w$ at distance at most $\ell/2-1$ from both $v_{i}$ and $v_{j}$. By the definition of $N_{i}$ and $N_{j}$, $w\in N_{i}\cap N_{j}$, which proves our statement. ∎
Now, Theorem 1 implies that there is a absolute constant $T$ such that there exists a point set $Z=\{z_{1},z_{2},\dots,z_{T}\}$ hitting each member of $\mathcal{N}$.
Note that we may take the points of $Z$ to be vertices of $G$, as the sets in $\mathcal{N}$ are unions of neighborhoods of vertices and edges of $G$, and if a set $N_{i}$ contains a point in the neighborhood of an edge $e=uv$, then $N_{i}$ contains both $u$ and $v$. As a consequence, we have a set of $T$ vertices such that for each vertex $v_{i}\in C$, there is an element of $Z$ at distance at most $\ell/2-1$ from $v_{i}$ in $G$.
Inspired by Axenovich, Kießle and Sagdeev [3], we define a partitioning of $C$ into sets $C_{1},C_{2},\dots,C_{T}$ such that $v_{i}\in C_{j}$ if and only if $d(v_{i},z_{j})\leq d(v_{i},z_{j^{\prime}})$ for each $j^{\prime}\neq j$, and $d(v_{i},z_{j})<d(v_{i},z_{j^{\prime}})$ for each $j^{\prime}<j$. In other words, we assign each $v_{i}$ to the minimum index vertex in $Z$ which is closest to it.
Next, we present two statements about the distributions of the sets $C_{i}$, which help us bound the number of vertices on $C$.
The following observation was used by Axenovich, Kießle and Sagdeev [3] as well, we include its proof for completeness:
Lemma 13 ([3]).
If $v_{i+1},\dots,v_{i+j}\in C_{k}$ are consecutive vertices of $C$ in the same partition class, then $j\leq\ell-1$.
Proof.
Since for each $h$, the union of the shortest paths from $v_{h}$ to $z_{k}$ and from $v_{h+1}$ to $z_{k}$, and the edge $v_{h}v_{h+1}$ forms a closed walk of length at most $2(\ell/2-1)+1=\ell-1$, this walk has to be trivial, and thus $|(d(v_{h},z_{k})-d(v_{h+1},z_{k})|=1$ must hold.
Additionally, if there was an index $h$ such that $d(v_{h-1},z_{k})+1=d(v_{h},z_{k})=d(v_{h+1},z_{k})+1$, then there would be a non-trivial closed walk of length at most $2(\ell/2-1)=\ell-2$ through $v_{h}$ and $z_{k}$, which is a contradiction.
As a corollary, the function $h\mapsto-d(v_{h},z_{k})$ is strictly unimodal. Since $0\leq d(v_{h},z_{k})\leq\ell/2-1$, it can take at most $2(\ell/2-1)+1=\ell-1$ different values, proving our result. ∎
Lemma 14.
There cannot exist indices $h_{1}<h_{2}<h_{3}<h_{4}$, for which $v_{h_{1}},v_{h_{3}}\in C_{i}$ and $v_{h_{2}},v_{h_{4}}\in C_{j}$ for some $i\neq j$.
Proof.
Assume without loss of generality that $i<j$. Let $P_{a}$ be a shortest path between $z_{i}$ and $v_{h_{a}}$ for $a\in\{1,3\}$, and similarly, let $P_{b}$ be a shortest path between $z_{j}$ and $v_{h_{b}}$ for $b\in\{2,4\}$. As $h_{1}<h_{2}<h_{3}<h_{4}$, there must be an $a\in\{1,3\}$ and $b\in\{2,4\}$ for which $P_{a}$ and $P_{b}$ intersect in a vertex $w$. If $d(w,z_{i})\leq d(w,z_{j})$, then $d(v_{h_{b}},z_{i})\leq d(v_{h_{b}},z_{j})$ which contradicts the definition of the sets $C_{i},C_{j}$. Similarly, if $d(w,z_{i})>d(w,z_{j})$, then $v_{h_{a}}$ should be in $C_{j}$ rather than $C_{i}$. ∎
We use a theorem of Davenport and Schinzel [10]:
Theorem 15 ([10]).
If we color the integers $\{1,2,\dots,n\}$ with $t$ colors such that neighboring indices have different colors, and there are no indices $h_{1}<h_{2}<h_{3}<h_{4}$ and colors $c_{i},c_{j}$ such that $h_{1},h_{3}$ have color $c_{i}$ and $h_{2},h_{4}$ have color $c_{j}$, then $n\leq 2t-1$.
Color vertices of $C_{i}$ with color $c_{i}$ for each $i$, and contract each monochromatic interval, thus getting an interval colored with $T$ colors which satisfies the conditions of Lemma 15. Each contracted interval had length at most $\ell-1$, therefore the length of $C$ must satisfy $|C|\leq(2T-1)(\ell-1)$, which completes the proof of Theorem 7 and thus of Theorem 2. ∎
4 Face lengths on other surfaces
In this section, we show the following analogue of Theorem 2 for graphs drawn on orientable surfaces:
[…]
The proof will proceed similarly, with the main difference being the usage of a new result of Raman and Singh [21] on supports for subgraph systems on surfaces.
Instead of considering blow-ups of subgraphs and using their non-piercing property, we will consider the subgraphs themselves, with the non-piercing property exchanged to cross-freeness, which was introduced by Raman and Singh [21]:
Definition 16.
For a fixed graph $H$ drawn on a surface and a pair of subgraphs $H_{1},H_{2}$, the reduced graph $R(H_{1},H_{2})$ is the graph obtained by contracting all edges, both of whose end-points are in $H_{1}\cap H_{2}$.
Consider two subgraphs $H_{1},H_{2}$ and a vertex $v$, and let $\tilde{v}$ be the image of $v$ in $R(H_{1},H_{2})$. Subgraphs $H_{1}$ and $H_{2}$ are said to be crossing at a vertex $v$ if there are edges $e_{1},e_{2},e_{3},e_{4}$ in $R(H_{1},H_{2})$ incident to $\tilde{v}$ in this circular order, with $e_{1},e_{3}\in H_{1}\setminus H_{2}$ and $e_{2},e_{4}\in H_{2}\setminus H_{1}$.
A set of subgraphs $\mathcal{H}=\{H_{1},H_{2},\dots,H_{k}\}$ without crossing pairs is called a cross-free (or non-crossing) graph system.
We first show the following analogue of Theorem 1:
Definition 17.
An $r$-neighbourhood around a vertex $u$ in a graph $G$ is the subgraph of $G$ induced by the vertices at distance at most $r$ from $u$ in $G$.
Lemma 18.
Let $G$ be a graph drawn on a surface with genus $g$ and girth larger than $2r+1$, and let $\mathcal{H}=\{H_{1},H_{2},\dots,H_{k}\}$ be a pairwise intersecting family of its $r$-neighbourhoods. Then $\mathcal{H}$ is a cross-free family and it can be pierced with at most $C(g)$ points for some $C(g)$ constant only depending on $g$.
Proof.
We again show a bound for the dual VC-dimension and show a $(p,q)$-property for $\mathcal{H}$, and we conclude using Matoušek’s theorem.
Observation 19.
$H_{i}$ is a tree for each $i$.
Proof.
If $H_{i}$ contained a cycle, it would have length at most $2r+1$, contradicting the girth constraint. ∎
Lemma 20.
The dual VC-dimension of $\mathcal{H}$ is less than $\frac{7+\sqrt{49+48g}}{2}$.
Proof.
Our proof starts similarly to Lemma 4. Suppose, for a contradiction, that we have subgraphs $H_{1},\dots,H_{\varphi}$ for $\varphi=\lceil\frac{7+\sqrt{49+48g}}{2}\rceil$ exhibiting this dimension, and choose points $x_{i}\in H_{i}\setminus\bigcup_{k\neq i}H_{k}$, $x_{i,j}\in\left(H_{i}\cap H_{j}\right)\setminus\bigcup_{k\neq i,j}H_{k}$. Take simple curves $\gamma_{i,j}$ to be paths from $x_{i}$ to $x_{i,j}$ inside $H_{i}$, and we again get an embedding of $K_{\varphi}$ on the surface. More precisely, to get a drawing where any two edges meet a finite number of times, we need to make a small perturbation to make the curves $\gamma_{i,j}$ edge-internally disjoint, i.e., any two curves can intersect only at a small neighborhood of a vertex. Moreover, we can also assume that for any $i$ the curves $\{\gamma_{i,j}\}_{j}$ are pairwise disjoint (apart from all starting from $x_{i}$), as otherwise we could “uncross” such intersections. As the minimal genus of a surface into which $K_{\varphi}$ can be embedded is $\lceil\frac{(\varphi-3)(\varphi-4)}{12}\rceil>g$ [22], there will be two independent intersecting edges, without loss of generality, $\gamma_{1,2}$ and $\gamma_{3,4}$.
Suppose that $H_{1}$ and $H_{3}$ are neighbourhoods around $y_{1}$ and $y_{3}$, respectively, and let $z$ be a vertex in the small vicinity of an intersection point of $\gamma_{1,2}$ and $\gamma_{3,4}$.
By Observation 19, $H_{1}$ is a tree. Therefore, since the vertices $x_{1},z,x_{1,2}$ lie on a path in this order, either $d(x_{1},y_{1})=d(x_{1},z)+d(z,y_{1})$ or $d(x_{1,2},y_{1})=d(x_{1,2},z)+d(z,y_{1})$. Similarly, either $d(x_{3},y_{3})=d(x_{3},z)+d(z,y_{3})$ or $d(x_{3,4},y_{3})=d(x_{3,4},z)+d(z,y_{3})$.
First, assume that the first equality holds both times. By symmetry, we may assume $d(x_{1},z)\geq d(x_{3},z)$. Then,
$d(x_{3},y_{1})\leq d(x_{3},z)+d(z,y_{1})\leq d(x_{1},z)+d(z,y_{1})=d(x_{1},y_{1}),$
which implies $x_{3}\in H_{1}$ as $x_{1}\in H_{1}$, a contradiction since $x_{3}\in H_{3}\setminus H_{1}$.
We also have $x_{1,2}\in H_{1}\setminus H_{3}$ and $x_{3,4}\in H_{3}\setminus H_{1}$, so it does not matter which of the equalities hold, we could repeat the same proof with $x_{1,2}$ instead of $x_{1}$ and/or $x_{3,4}$ instead of $x_{3}$. ∎
Proposition 21.
The subgraph-system $\mathcal{H}=\{H_{1},H_{2},\dots,H_{k}\}$ is cross-free.
Proof.
Assume that two subgraphs, without loss of generality, $H_{1}$ and $H_{2}$, cross at a vertex $z$. This implies the existence of vertices $x_{1},x_{2}\in H_{1}\setminus H_{2}$ and $x^{\prime}_{1},x^{\prime}_{2}\in H_{2}\setminus H_{1}$ such that $z$ lies on the unique path between $x_{1}$ and $x_{2}$ in $H_{1}$, and it also lies on the unique path between $x_{1}^{\prime}$ and $x_{2}^{\prime}$ in $H_{2}$, and we reach a contradiction the same way as in the proof of the previous lemma. ∎
Recall that for a set-system $\mathcal{F}$ on a set of elements $P$, we define the dual intersection hypergraph $\mathcal{H}(\mathcal{F},P)$ such that the vertices correspond to members of $\mathcal{F}$, while hyperedges correspond to elements of $P$ such that for every $p\in P$ the vertex set $H_{p}=\{F\in\mathcal{F}:p\in F\}$ forms a hyperedge. The Delaunay graph $\mathcal{D(\mathcal{F})}$ is the subgraph of $\mathcal{H}(\mathcal{F},P)$ containing only the hyperedges with exactly two vertices—so each edge $F_{1}F_{2}$ is witnessed by some exclusive intersection point $x\in F_{1}\cap F_{2}$ such that no $F\in\mathcal{F}\setminus\{F_{1},F_{2}\}$ contains $x$.
Raman and Singh [21] have shown that for any non-crossing graph system $\mathcal{F}$ on a graph $G_{\mathcal{F}}$ of genus $g$, there exists a dual support of $\mathcal{F}$ on genus at most $g$, that is, there exists a graph $G_{\mathcal{F}}^{\prime}$ of genus $g$ such that for each $p\in V(G_{\mathcal{F}})$, $H_{p}$ is a connected subgraph of $G_{\mathcal{F}}^{\prime}$.
We have the following consequence for the Delaunay graph $\mathcal{D(\mathcal{H})}$, where each exclusive intersection point corresponds to an edge.
Corollary 22 (of Raman and Singh [21] ).
For any non-crossing graph system $\mathcal{F}$ on a graph $G_{\mathcal{F}}$ of genus $g$, the Delaunay graph $\mathcal{D}(\mathcal{\mathcal{F}})$ can be drawn on a surface with genus $g$, therefore it can have at most 3|F|6+6g3|F|-6+6g edges.
Lemma 23.
If $\mathcal{F}$ is a family non-crossing subgraphs with the $(\nu+1,2)$-property, then $\mathcal{F}$ also has the $(p,q)$-property for any $p>3e(\nu+1)\nu q+2gq+1$ for any $q\geq 2$, where $e=2.71$... is Euler’s number.
Proof.
The proof analogous to that of Lemma 6, with the only difference being a different upper bound on the number of edges in the Delaunay graph.
Fix some $q\geq 2$ and $p>3e(\nu+1)\nu q+1$, and suppose for a contradiction that there exists a subfamily $\mathcal{G}\subset\mathcal{F}$ of size $p$ without a common intersection. Delete each set from $\mathcal{G}$ with probability $1-\frac{1}{q}$ to obtain the subsubfamily $\mathcal{H}$. The average number of remaining sets is $\frac{p}{q}$, so in expectation the Delaunay graph $\mathcal{D(\mathcal{H})}$ has this many vertices. The probability that two different intersecting sets of $\mathcal{G}$ span an edge of $\mathcal{D(\mathcal{H})}$ is at least $\frac{1}{q^{2}}\left(1-\frac{1}{q}\right)^{q-3}>\frac{1}{e}\frac{1}{q^{2}}$, as any point in the intersection of two sets is contained in at most q3q-3 other sets; if these are all deleted, then we get a Delaunay edge. As from any $\nu+1$ sets in $\mathcal{G}$, there are two that intersect, the number of intersecting set pairs in $\mathcal{G}$ is at least $\frac{\binom{p}{\nu+1}}{\binom{p-2}{\nu-1}}=\frac{p(p-1)}{(\nu+1)\nu}$. Consequently, the expected number of edges is at least $\frac{1}{e}\frac{p(p-1)}{(\nu+1)\nu q^{2}}$. As $\mathcal{G}$ and all its subsystems are non-crossing, by Corollary 22, we get the following inequality for the expected number of edges after deletion: $\frac{1}{e}\frac{p(p-1)}{(\nu+1)\nu q^{2}}\leq 3\frac{p}{q}+6g$. Reordering, we get $p(p-(3e(\nu+1)\nu q+1))\leq 6eg(\nu+1)\nu q^{2}$, which contradicts $p>3e(\nu+1)\nu q+2gq+1$. ∎
As in the proof of Theorem 1, we conclude using Theorem 3. ∎
Proof of Theorem 2.
Fix a graph $G$ with girth at least $\ell$ drawn on a surface with genus $g$, and fix a face of $G$ with facial cycle $C=v_{1}v_{2}\dots v_{m}$. Similarly as in the proof of Theorem 2, we may assume that $\ell$ is even.
For every vertex $v_{i}\in C$, define its $\left(\frac{\ell}{2}-1\right)$-neighbourhood $N_{i}=\{\ v\in V(G):d(v,v_{i})\leq\frac{\ell}{2}-1\}\cup\{e\in E(G):d(e,v_{i})\leq\frac{\ell}{2}-1\}$.
Lemma 24.
The subgraph-system $\mathcal{N}$ has the $(2,2)$-property, i.e., it is pairwise intersecting.
Proof.
Take two arbitrary vertices $v_{i},v_{j}\in C$. By Observation 8, $d(v_{i},v_{j})\leq\ell-2$. Since $\ell$ is even, this implies that there is a vertex $w$ at distance at most $\ell/2-1$ from both $v_{i}$ and $v_{j}$. By the definition of $N_{i}$ and $N_{j}$, $w\in N_{i}\cap N_{j}$, which proves our statement. ∎
Now, we can use Lemma 18 with $r=\frac{\ell}{2}-1$, getting $C(g)$ points $z_{1},z_{2},\dots,z_{C(g)}$ covering all neighbourhoods $N_{i}$.
Again, we define a partioning of the vertices of the cycle such that $v_{i}\in C_{j}$ if and only if $d(v_{i},z_{j})\leq d(v_{i},z_{j^{\prime}})$ for each $j^{\prime}\neq j$, and $d(v_{i},z_{j})<d(v_{i},z_{j^{\prime}})$ for each $j^{\prime}<j$.
As Lemma 13 does not use any topological properties of the surface, it still holds here. In contrast, we need to show a substitute for Lemma 14:
Lemma 25.
There cannot exist indices $i_{1}<j_{1}<i_{2}<j_{2}<\dots<j_{b}$, for which $v_{i_{1}},v_{i_{2}},\dots,v_{i_{b}}\in C_{i}$ and $v_{j_{1}},v_{j_{2}},\dots,v_{j_{b}}\in C_{j}$ for some $i\neq j$, where $b=4g+3$.
Proof.
Assume that there exist two such sets $C_{i},C_{j}$. We will show that this implies the the existence of a proper embedding of $K_{3,b}$ on the vertex set $\{z_{i},z_{j},z\}\cup\{v_{i_{i}},v_{i_{2}},\dots,v_{i_{b}}\}$, where $z$ is an arbitrary point of the empty face.
For $h\in[b]$, let the edge $v_{i_{h}}z_{i}$ of the embedding be the shortest path between $v_{i_{h}}$ and $z_{i}$ in the graph, and let the edge $v_{i_{h}}z_{j}$ of the embedding be concatenation of the shortest path between $v_{j_{h}}$ and $z_{i}$, and the path between $v_{i_{h}}$ and $v_{j_{h}}$ on $C$. Note that each path from $z_{i}$ is disjoint from each path from $z_{j}$, and the paths from $z_{i}$ (respectively, from $z_{j}$) can be made internally pairwise disjoint with a small perturbation. Finally, $z$ can be connected to vertices $v_{i_{h}}$ with arbitrary, pairwise non-intersecting curves.
It is easy to see that these curves yield a proper embedding of a $K_{3,b}$ graph on the surface, which implies $b\leq 4g+2$, as the genus of the complete bipartite graph $K_{p,q}$ is $\lceil\frac{(p-2)(q-2)}{4}\rceil$ [23].
Here, we use the general form of the Davenport-Schinzel theorem [10]: Let $\lambda_{b}(t)=t\cdot 2^{\frac{1}{s!}\alpha(t)^{s}+O\left(\alpha(t)^{s-1}\right)}$, where $s=\lfloor\frac{b-2}{2}\rfloor$ and $\alpha(t)$ is the inverse Ackermann function.
Theorem 26 ([10]).
If we color the integers $\{1,2,\dots,n\}$ with $t$ colors such that neighboring indices have different colors, and there are no indices $h_{1}<h_{2}<\dots<h_{2b-1}<h_{2b}$ and colors $c_{i},c_{j}$ such that $h_{1},h_{3},\dots,h_{2b-1}$ have color $c_{i}$ and $h_{2},h_{4},\dots,h_{2b}$ have color $c_{j}$, then $n\leq\lambda_{b}(t)$.
Again, we color vertices of $C_{i}$ with color $c_{i}$ for each $1\leq i\leq C(g)$, and contract each monochromatic interval, thus getting an interval colored with $C(g)$ colors satisfying the conditions of Lemma 25. By Lemma 13, each contracted interval had length at most $\ell-1$, therefore $|C|\leq\lambda_{4g+3}(C(g))(\ell-1)$. ∎
Concluding remarks and open questions
There are several questions left open, from which we would like to highlight a few.
Question 1.
Is it true that every non-piercing pairwise intersecting family can be pierced with 4 points? What about pseudo-disks?
If yes, the method of our proof for Theorem 2 would give an upper bound of $7\ell-7$ for $f_{max}(\ell)$, improving the current best bound $8\ell-13$ of Axenovich, Kießle, Sagdeev, and Zhukovskii [3].
Question 2.
What is the exact constant for $f_{max}(\ell)?$ Is it possible to improve the constructions in [3]? What is the exact characterization of maximal $\mathcal{C}_{<\ell}$-free graphs?
Acknowledgement
We would like to thank Arsenii Sagdeev for discussion, in particular, to pointing out that a positive answer to Question 1 would give a better upper bound for $f_{max}(\ell)$. He also pointed out that the methods of their paper and ours might be combined to obtain an improved upper bound, but we leave this for future research.